OpenStudy (anonymous):

Mason walked 1/3 mile before lunch and 5/12 mile after lunch. How many miles in all?

OpenStudy (anonymous):

3/4

OpenStudy (johnweldon1993):

\[\large \frac{1}{3} + \frac{5}{12}\] find the least common denominator ...

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

=3/4

OpenStudy (anonymous):

oh thanks!

OpenStudy (anonymous):

your welcome BYE